OpenStudy (anonymous):

Solve x2 + 10x + 24 = 0 by completing the square. Show your work

OpenStudy (hoblos):

x2 + 10x + 24 = 0 x2 + 10x + 24 +1-1= 0 x2 + 10x + 25 -1 = 0 x2 + 10x + 25 = 1 (x+5)^2=1 x+5 = ±√1 = ±1 so x=-6 or x=-4